If father age is four times than his son and after 24 years, father age is twice as his son. What is their present age?
Answer by Cromlix(4381)

Hi there,
If father age is four times than his son
F = 4S
after 24 years, father age is twice as his son
F + 24 = 2(S + 24)
Replace F with 4S from first equation.
4S + 24 = 2(S + 24)
4S + 24 = 2S + 48
Collect like terms
4S - 2S = 48 - 24
2S = 24
S = 12
Son = 12 years old
Father = 48 years old
Hope this helps :-)